The width of a rectangle is 5 more than one third it's length. Of the perimeter is 34, what are the dimensions of the rectangle?
Answer by stanbon(75887)

The width of a rectangle is 5 more than one third it's length. Of the perimeter is 34, what are the dimensions of the rectangle?
------
Perimeter = 2(length + width)
----
34 = 2(L + (1/3)L +5)
---
17 = (4/3)L + 5
(4/3)L = 12
Length = 9
------------------------------
width = (1/3)9 + 5
width = 8
==================================
Cheers,
Stan H.
==================================